OpenStudy (anonymous):

Solve the inequality 1/3x+3/4<1/2x and describe the graph of the solutions. Show all work

OpenStudy (anonymous):

Our goal is to isolate the x to one side of the inequality. We kind of treat the inequality sign like an equal sign and solve for x The first step is to bring all the x's to one side of the inequality So we would begin by subtracting a 1/3x from both sides \(\large \frac{1}{3}x- \frac{1}{3}x+ \frac{3}{4}<\frac{1}{2}x-\frac{1}{3}x\) \( \large \frac{3}{4}<\frac{1}{2}x-\frac{1}{3}x\) \( \large \frac{3}{4}<\frac{1}{6}x\)

OpenStudy (anonymous):

Ya with me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So now in order to isolate the x we multiply both sides by 6 \( \large \frac{3}{4}*\small{6}< \small{6}*\large\frac{1}{6}x \) \( \large \frac{18}{4}<x\) \( \large 4.5<x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Any questions so far and can you explain to me what x >4.5 means?

OpenStudy (anonymous):

x is greater than 4.5?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whooopsssss I meant So basically in order for this inequality to be true x must be greater than 4.5

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):